1)100 mg:1,000 mg::1 mL:X mL. Determine the value of X.

2) 100 mg:200 mg::1 mL:X mL. Determine the value of X.

3) Emily weights 54 lbs. If the doctor orders a medication at 20 mg/kg/dose with two doses per day, how much would she receive for each dose?

4) The order reads give 200 mg of amoxicillin. The label reads 100 mg/2ml. How many milliliters will you give?

5) If a drug is ordered at 10 mg/kg/day, how much would a 20 lb patient need each day?

6) A mother calls your office and says that she does not know how many teaspoons to give her child becasue the directions on the medication bottle read "Give 15 ml". How many teaspoons does that equal?

7) A drug label indicates you should add 3.4 ml diluent to powder containing one gram of antibiotic. The resulting solution will provide 250 mg of medication per ml. The order reads 300 mg. How much would you give the patient?

8) D = 750 H = 250 Q = 2. Using the formula method, calculate the proper dose
Answer #1

Solution 1: It is given 100 mg present in 1ml

100mg is present in 1ml

Therefore, 1000mg will be present in 1000/100

= 10ml.

Solution 2:  

100mg is present in 1ml

Therefore, 200mg will be present in = 200/100

= 2ml

Solution 3:

Weight of the patient = 54 lbs.

So, weight of the patient in kilogram= 54/2.205 = 24.494 kilogram.

Ordered drug = 20mg/kg/dose.

So, 24.494 × 20 = 485.88 mg.

Solution 4:

Ordered drug = 200mg

Available drug = 100mg / 2ml

So, 100 mg of the drug is present in 2ml

Therefore, 200mg of drug will be present in = 200 ×2 = 400

400÷ 100 = 4ml.